Minutes of the 14th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson Minutes taker Mark Dennehy Present Steven Watterson (Captain),

Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Conor McDermottroe (Public Relations Officer), Peter Byrne,

Absent(with apologies) Absent Alan Smyth (Chief Range Officer),

Deirdre Ninaber (Ladies’ Coordinator)

Date 20 November 2001 1945 – 2130 Distribution DURC Committee

Overview of topics

1 Preliminary discussion 1

2 Reports 1 2.1 Secretarys Report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2.2 Treasurers Report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 A&E Report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Committee Members 2

4 RO Dinner 2

5 10m Open 3

6 AOB 3 6.1 Regular Committee Time . . . . . . . . . . . . . . . . . . . . . . . . . 3

1 Preliminary discussion

6.2 World University Championships . . . . . . . . . . . . . . . . . . . . . 3 6.3 Grudge Match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1 Preliminary discussion

Before the meeting proper took place, John and Mark discussed the prizes for the DURC 10m Open. John was worried that Marks budget of £172 was over the preagreed £150 limit. Mark replied that the 172 was likely to be too high, but working from catalog prices did that and we usually got a good deal from Tom Watkins in Frederick St. Mark also said that the priority was to have a good open and not to make a profit - John replied that it would be impossible to absorb a loss from the Open so caution was necessary.

2 Reports

2.1 Secretarys Report

Mark reported on the correspondance, all rejections of sponsorship requests from various companies for the DURC Open.

2.2 Treasurers Report

Ben reported that the Capital Grant request had gone in, that the NonCapital grant request had gone in with the bill for repairing Air Rifle V, Competitions for the next few months, and supplies for lead removal (goggles and masks). Ben also reports at at 402 members, we are now the largest club in college. Ben reports that Simon Kennedy has posted us a cheque for his contribution to the DURC Open.

2.3 A&E Report

John reports lots of ammunition and pellets are in front arch. The Walther bolt was found on the ground over the weekend. The culprit is unknown, and noone is happy with the treatment of the bolt. The ammunition count is out. John will correct the problem. However, John is annoyed that it went for three days before being reported, and he will leave his mobile number in the locker for immediate notification of any future problems. Ben leaves

temporarily

Peter said that he had noticed that there wasn’t a lot of T22 in front arch - at last count, 230 rounds or so. John replied that the rate of ammunition usage coupled with the low limit of ammunition we are allowed store in Front Arch and Pat Moray’s not transferring ammunition except on Thursdays is causing problems in supply. Any ROs seeing the level is low should put in a request for more ammunition to Pat’s office with the provided forms.

Steven said that the pellets were running out faster than expected, with almost a full sleeve already shot. We need to watch the levels of pellets more carefully. John replied that at most only one-third to one-half of the sleeve could possibly have allready been

3 Committee Members

shot and people were probably just opening new tins a lot and then underestimating the amount of pellets remaining. Steven suggested finding all the less-than-full tins and consilidating them to full tins.

3 Committee Members

Steven thought the wording of the agenda - “Formal dismissal of Jason Sibley” was a bit curt. Mark replied that there was no negative meaning behind it, that that was just the name of the procedure.

The committee then voted unanimously to dismiss Jason from committee and elect Peter Byrne formally to committee, without a preselected portfolio.

4 RO Dinner

Mark reported a possible conflict with the WTSC dinner on the 15th. It was decided that in order for the dinner to go ahead, it had to be scheduled and just run as opposed to trying to appease everyone’s schedule.

The location was discussed. Verona’s was dismissed as expensive, LaLuna as being too small, Luigi Malones was dismissed when John reported a bad dinner there (though Mark reported eating there with no problems), and Tante Zoe’s was dismissed when John reported not liking their menu. Freemans was agreed upon as the location.

Numbers were talked about, and it was decided that 17 was a reasonable estimate. Mark is to book Freemans for this many people on the 15th, contingent on an a la carte menu being available as opposed to a set menu.

A deposit was agreed on at £10 per person by December 1st. Ben returns

5 10m Open

Alan is to draw up a new rota for ROing at the open as the current one makes no sense. Keith will score on sunday and Steven on saturday.

Transport was discussed. Peter will have his car on Sunday and possibly Saturday. Gear will be hauled out on the Friday evening. Mark will have his car for the full weekend.

Targets are not due until Friday. We may have to use NSRA targets.

Handicap schemes for the Team Shoot were discussed. This proved to be difficult to agree on, but eventually a class-based handicap was chosen, with points added to each team member’s gunscore according to their class.

Flavours for cheesecake were decided on for mark to cook.

6 AOB

6 AOB

6.1 Regular Committee Time

Mark pointed out that a regular time was needed for committee meetings as at the moment noone can prepare adaquately and there is the possibility of committee members being assigned to RO duty during a committee meeting. Tuesday evenings were picked, starting at 1830. Alan is to redraft the RO roster for the rest of the year with this constraint. Ben will see if he can book the IAMS room for the meetings.

6.2 World University Championships

Ben’s email on this competition was reported. Peter wondered about qualification procedures and scores. Mark said that Pat O’Brien was looking at this from the NTSA. Steven warned all that we would get trounced in this competition due to the standard of the other competitiors, most of whom were able to train nearly full-time in properly funded conditions with good facilities and excellent coaching. John was not happy at this as he felt DUCAC would unfairly compare us to these others and we’d lose our good image.

6.3 Grudge Match

The topic of grudge matches came up, with Peter wondering about a grudge match with the Karate club. No-one seemed keen, Mark pointing out that every member of the Karate club could tell the story of how they broke their noses for the first time, and that we allready have a grudge match with the Fencing club and have had one with Security before. The Building’s Office team was also mentioned.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

6 AOB

Tasks agreed on at meeting